OpenStudy (ammarah):

A restaurant manager bought 20 packages of bagels. Some packages contained 6 bagels each, and the rest contained 12 bagels each. There were 168 bagels in all. How many packages of 12 bagels did the manager buy?

OpenStudy (ammarah):

@dumbcow

OpenStudy (zzr0ck3r):

x+y = 20 6x+12y = 168

OpenStudy (zzr0ck3r):

so 6x+12(20-x)=168 solve for x, and that is the amount of packages with 6

OpenStudy (wolfboy):

Well first 168/12 is 14 And we need 6 more packages so scince 6 is half of 12 one package of twelve is two of six. 14-2=12=+4=16 16-2=14+4=18-2=16+4=20. So he bought 8 packages of twelve. Someone please correct me if i messed up.

OpenStudy (wolfboy):

@zzr0ck3r Did I or am I in the clear?

OpenStudy (zzr0ck3r):

well my way leads to 6x+240-12x=168 -6x=-72 x=12 so he bought 12 packages of 6, since x+y=20 implies 12+y = 20 thus y = 8 so he bought 8 packages of 12. correct:))

OpenStudy (wolfboy):

Lol Thank you for checking

OpenStudy (zzr0ck3r):

your logic is correct, but I assume they wanted a system of equations..but maybe not.... all good:)